OpenStudy (anonymous):

x+2/x^2-4 times x-2/x^2+4

OpenStudy (anonymous):

\[(\frac{x+2}{x ^{2}})\times(\frac{x-2}{x ^{2}+4})\] right?

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

\[\frac{(x+2)(x-2)}{x ^{2}(x ^{2}+4)}\]

OpenStudy (anonymous):

can't reduce any more